OpenStudy (anonymous):

The surface area of a cube is 96 ft^2. What is the length of each edge? (Yes, I realize this is probably a very simple question, but I have been learning about quadratics and I've forgotten most of this stuff.)

OpenStudy (anonymous):

First you know a cube has 6 faces, so you'd have to find the surface area of one face. Once you find that, then you also know that a square has equal sides and the area of one of its faces is the product of these sides.

OpenStudy (e.cociuba):

surface area of a cube x = length of edge surface area of 1 face is "x squared" written (x^2) a cube has 6 surfaces, total surface area = 6(x^2) 96cm2 = 6(x^2) solve for x

OpenStudy (anonymous):

I got x=4, is this right?

OpenStudy (e.cociuba):

yup u got it!

OpenStudy (anonymous):

Awesome! Thank you for helping.

OpenStudy (e.cociuba):

np:)